# Planning behavior in a recurrent neural network that plays Sokoban

### v1

@@ Abstract
To predict how advanced neural networks generalize to novel situations, it is essential to understand how they reason. Guez et al. (2019, “An investigation of model-free planning”) trained a recurrent neural network (RNN) to play Sokoban with model-free reinforcement learning. They found that adding extra computation steps to the start of episodes at test time improves the RNN’s success rate. We further investigate this phenomenon, finding that it rapidly emerges early on in training and then slowly fades, but only for comparatively easier levels. The RNN also often takes redundant actions at episode starts, and these are reduced by adding extra computation steps. Our results suggest that the RNN learns to take time to think by ‘pacing’, despite the per-step penalties, indicating that training incentivizes planning capabilities. The small size (1.29M parameters) and interesting behavior of this model make it an excellent model organism for mechanistic interpretability.
Machine Learning, ICML
Figure 1: Top: Proportion of medium-difficulty validation levels solved vs. environment steps used in training. Curves show the RNN’s performance with a specific number of forced thinking steps at episode start, along with a ResNet baseline. Bottom: Estimated planning effect: the 8-steps minus the 0-steps curve. We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels, but decreases for medium levels.
@@ 1 Introduction
In many tasks, performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether giving a human more time to think before taking a chess move, or prompting a large language model to reason step by step (Kojima et al., 2022). Thus, understanding reasoning in such NNs is key to predicting their behavior in novel situations.
@@
Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment. So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning (Hubinger et al., 2019) – may internalize goals different from that given by the training objective, leading to harmful misgeneralization (Di Langosco et al., 2022; Shah et al., 2022).
This paper presents a 1.29M parameter recurrent neural network (RNN) that clearly benefits from extra time to reason. Following Guez et al. (2019), we train a convolutional LSTM to play Sokoban, a challenging puzzle game that remains a benchmark for planning algorithms (Peters et al., 2023) and reinforcement learning (Chung et al., 2024).
We believe understanding this RNN is a useful first step towards understanding reasoning in sophisticated neural networks. The network is complex enough to be interesting: it reasons in a challenging environment and benefits from extra thinking time. At the same time, it is simple enough to be tractable to reverse engineer with current mechanistic interpretability techniques, owing to its small size and relatively straightforward behavior.
We make three key contributions. First, we replicate and open-source an RNN whose performance improves with thinking time (Guez et al., 2019). Second, we analyze its training dynamics, finding that the thinking-time behavior emerges early in training and is incentivized by the setup (Figure 1, Sections 3 and 4.1). Third, we conduct a detailed empirical analysis of the network’s behavior. We quantitatively investigate how increased thinking changes network behavior (Section 3), showing that it reduces myopic actions (Figure 5). Furthermore, we present case studies of planning behavior that suggest the RNN naturally takes time to think (Section 4), which may explain the reduced planning effect on medium levels (Figure 1).
2 Training the test subject
@@
We closely follow the setup from Guez et al. (2019), using the IMPALA (Espeholt et al., 2018) reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture, as well as a ResNet. We open-source both the trained networks [^1] https://github.com/AlignmentResearch/learned-planner and training code [^2] https://github.com/AlignmentResearch/train-learned-planner.
@@
Sokoban is a grid puzzle game with walls, floors, movable boxes, and target tiles. The player’s goal is to push all boxes onto target tiles while navigating walls. We use the Boxoban dataset (Guez et al., 2018), consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets. The edge tiles are always walls, so the playable area is $8\times 8$. Boxoban separates levels into train, validation and test sets, with three difficulty levels: unfiltered, medium, and hard. Guez et al. (2019) generated these by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels. In this paper, we use unfiltered-train (900k levels) to train networks. To evaluate them, we use unfiltered-test (1k levels) [^3] We prefer unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al. (2019)., medium-validation (50k levels), and hard ($\sim$3.4k levels), which do not overlap.

### v2

# Planning in a recurrent neural network that plays Sokoban

@@ Abstract
How a neural network (NN) generalizes to novel situations depends on whether it has learned to select actions heuristically or via a planning process. Guez et al. [2019, “An investigation of model-free planning”] found that a recurrent NN (RNN) trained to play Sokoban appears to plan, with extra computation steps improving the RNN’s success rate. We replicate and expand on their behavioral analysis, finding the RNN learns to give itself extra computation steps in complex situations by “pacing” in cycles. Moreover, we train linear probes that predict the future actions taken by the network and find that intervening on the hidden state using these probes controls the agent’s subsequent actions. Leveraging these insights, we perform model surgery, enabling the convolutional NN to generalize beyond its $10\times 10$ architectural limit to arbitrarily sized inputs. The resulting model solves challenging, highly off-distribution levels. We open-source our model and code, and believe the neural network’s small size (1.29M parameters) makes it an excellent model organism to deepen our understanding of learned planning.
[…]
Figure 1: Top left: Proportion of medium-difficulty validation levels solved vs. environment steps used in training. Curves show the RNN’s performance with a specific number of thinking steps (forced no-op actions) at episode start, along with a ResNet baseline. Bottom left: Estimated planning effect: the 8-steps minus the 0-steps curve. We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels (red), but decreases for medium levels (blue). Right: Linear probe predictions (arrows) for the direction the box will move in. Opacity is proportional to the number of steps in the episode in which the probe predicted a particular arrow. Correct predictions are in green and incorrect in red, the red predictions show other plausible plans for Box 4. This probe causally affects the actions of the agent, as described in Section 4. Boxes are numbered in the order in which they arrive at their final target.
@@ 1 Introduction
In many tasks, the performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether by giving a human more time to think before taking a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
@@
Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment. So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to harmful misgeneralization [Di Langosco et al., 2022, Shah et al., 2022].
Within goal-oriented reasoning, this work distinguishes between plans and search algorithms. A plan is a sequence of (perhaps abstract) actions which an agent intends to take. A search algorithm is an algorithm that considers many plans, evaluates them according to their predicted outcomes, and picks the plan evaluated as best. Search algorithms perform better the more plans they can evaluate and, with enough compute, generalize very well to novel problem instances [Russell and Norvig, 2009].
This work investigates a recurrent neural network (RNN) that plays Sokoban [Guez et al., 2019], a challenging puzzle game that remains a benchmark for planning algorithms [Peters et al., 2023] and reinforcement learning [Chung et al., 2024]. Using linear location-wise probes from concurrent work [Bush et al., 2024], we show that its activations represent a plan in a way which causes the RNN’s actions.
The behavior of the RNN suggests that it also performs search. We find that the RNN sometimes takes actions which return to previous states of the environment. This overwhelmingly happens at the beginning of episodes and can be substituted by no-ops, suggesting the RNN deliberately goes in cycles to get more time to think. Giving the RNN time to think with no-ops [Guez et al., 2019] disproportionately helps with levels that are difficult (Figure 1), long, or that require waiting many steps before the first reward (Figure 3).
Using insights from the plan probes and inspecting activations, we perform model surgery to uncover latent generalization capabilities of the RNN. The RNN is trained only on $10\times 10$ levels with 4 boxes (and its output MLP can only handle $10\times 10$ levels), but we are able to make it play on larger, nontrivial levels with more boxes (Figure 2).
Altogether, evidence from probes and behavior show that the RNN causally represents a plan. The evidence highly suggests that the RNN implements an online search algorithm which scales with computation, although we cannot rule out multi-step heuristics.
Consequently, we believe this RNN is an excellent model organism for understanding planning in sophisticated neural networks. Its behavior is complex enough to be interesting and its size (1.29M parameters) is small enough to be tractable to reverse engineer. This work shows that supervised probes and sparse autoencoders [Huben et al., 2023, Bricken et al., 2023, Gao et al., 2024] are not enough to fully do so, but it is likely that other mechanistic interpretability techniques such as circuit analysis [Olah et al., 2020, Geiger et al., 2021, Conmy et al., 2023] will reveal more about the algorithm which produces plans.
Figure 2: The $\text{DRC}(3,3)$ is limited to $10\times 10$ grids by a fixed-dimension MLP. With model surgery, the ConvLSTM backbone generalizes to much more challenging settings than the $10\times 10$ training set grids. Left: XSokoban-31, which the $\text{DRC}(3,3)$ solves with 0 thinking steps. Right: XSokoban-29, which the $\text{DRC}(3,3)$ solves but only after 128 thinking steps. More examples in Appendix B.
2 Setting up the test subject
@@
We train an agent closely following the setup from Guez et al. [2019], using the IMPALA [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture. We also train a ResNet baseline. We open-source both the trained networks  [^1] https://github.com/AlignmentResearch/learned-planner and training code  [^2] https://github.com/AlignmentResearch/train-learned-planner . For more details about architectures and training, see Appendix A.
$\text{DRC}(D,N)$ architecture.
Most of this paper focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019]. The core of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step. The output of the last ConvLSTM layer (the $D$th layer) is fed to the input at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to output the next action. We take $D=3,N=3$.
A linear combination of the mean- and max-pooled activations of the ConvLSTM is also fed to the next step, thus letting it communicate quickly across the receptive field (known as pool-and-inject). An encoder block consisting of two $4\times 4$ convolutions processes the input, which is also fed to each ConvLSTM layer. At the end, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
@@
Sokoban is a grid puzzle game with walls, floors, movable boxes, and target tiles. The player’s goal is to push all boxes onto target tiles while navigating walls. We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets. The edge tiles are always walls, so the playable area is $8\times 8$. Boxoban separates levels into train, validation and test sets, with three difficulty levels: unfiltered, medium, and hard. Guez et al. [2019] generated these sets by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels. In this paper, we use the unfiltered-train (900k levels) set to train networks. To evaluate them, we use the unfiltered-test (1k levels) [^3] We use unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al. [2019]., medium-validation (50k levels), and hard ($\sim$3.4k levels) sets, which do not overlap. To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix B).